40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Sun Nov 12 22:34:50 2017, Craig, awade, DailyProgress, NoiseBudget, Nov 12 Noisebudget PLL_OLG_SR785_Math_FMDevn_1kHz_PreampGain_2000_12-11-2017_195046_TF.pdfPLL_OLG_SR785_Math_FMDevn_1kHz_PreampGain_2000_12-11-2017_195046.tgz20171113_102854noiseBudget.pdf
    Reply  Mon Nov 13 15:45:19 2017, Craig, awade, HowTo, NoiseBudget, Nov 12 Noisebudget revised 20171113_151130noiseBudget.pdfPLLControlSignalSpectrum_Avg_30_FMDevn_1kHz_PreampGain_2000_Span_102p4kHz_12-11-2017_211211_Spectrum.pdf
       Reply  Mon Nov 13 18:06:32 2017, awade, Craig, DailyProgress, Schedule, Task list Nov 13, 2017 NorthLowerPeriscopeMirror_2017-11-13_16.48.15.jpgSouthLowerPeriscopeMirror_2017-11-13_16.49.35.jpg
          Reply  Mon Nov 13 21:53:25 2017, rana, DailyProgress, Schedule, Task list Nov 13, 2017 
Message ID: 1980     Entry time: Mon Nov 13 18:06:32 2017     In reply to: 1979     Reply to this: 1981
Author: awade, Craig 
Type: DailyProgress 
Category: Schedule 
Subject: Task list Nov 13, 2017 

Now that we have a noise spectra that touches the previous best, we should start making task lists to tackle the various noise floors to get to budget. I've drafted a list below as a starting point, edit this post with additions.

  • Investigate if glitches in FSS error signal come from the acromags, could be on of the binary switches or grounding coupling noise back into the interface box to the field box.  Check grounding is correct, monitor state of binary switches to see if they are near a threshold point, unplug and run in manual mode.
  • have a look at dust on critical mirrors in the final part of the beam path. Of the two LF noise shoulders in the noise budget plot (PSL:1979) the bottom one (≤ 100 Hz) is likely scatter.  Attached below are two pictures of the periscope mirrors under the green torch, there is a lot of dust on these (most likely foam dust).  I can see laser light scattering with the IR viewer, that means it pretty bad;
  • the second LF shoulder could be more scattering.  I haven't seen this extend far beyond 100-150 Hz as a broadband excitation for a setup with good optics mounting with 1" posts*, although you may see peaks from particular mounts.  This could be also FSS noise;
  • 636 Hz and 1040 Hz peaks seem to excited by tapping the table, these could be resonances of mounts (although its very high): do a simple tap or buzz test of a few mounts to see if we can pinpoint the offending element;
  • investigate breathing of north path error signal.  Seems to be on the order of 4 Hz;
  • setup Intensity Stabilization Servos (ISS) for north and south paths, check impact on BN spectrum; and
  • resolve mystery of why south FSS has error signal RMS of 5-10x that of north path (150 mVrms at the moment)

Future:

  • Prepare electronics for mode cleaner installation;
  • finish construction of BB EOM drivers for 36/7 MHz
  • check 36/7 MHz resonant detectors are actually in working condition and modify for new modulation frequencies
  • Window shields, attach sensors to new heat shields, electronics for in vacuum heaters/sensors.

Copy paste these check/cross marks for completed/dropped items: ___ ✓ ___ /___ ✗ ___

--- 

* at least in ridged 1" most mounted optical setups I've used before.

Quote:

REVISED BOKEH NOV 12 2017 NOISEBUDGET.

I realized that the Nov 12 Spectrum 4 Hz frequency comb an strong spikes at 256 and 512 Hz are not true noise artifacts, but merely poor measurement post processing.

In particular, when spectra are stitched together, there is sometimes some overlap in the frequency domain.  When this happens, the data taken at lower resolution is supposed to be thrown out.  I neglected to do this, but now have fixed the code in iris to do this.  awade calls this the Hierarchical Chucker.  

When I measured a spectrum with high span up to 102.4 kHz, the first point is at 0 Hz, and the second is at 256 Hz, and so on.  These points are inflated with all low frequency noise, and should be thrown away if we have spectra taken at higher resolution.  When one includes these points, we get the blue 'Nov 12 2017 Beat' line below.  When the points are properly hierarchically chucked, you get the gold 'Nov 13 2017 Beat' line.  These are the same measurement, just with some different post processing, so they ought to lie right on top of one another, which they do except for the line artifacts.

I have also included a plot of all the spectra so people can see why a comb can appear where it shouldn't if you don't throw out the low frequency data of your low resolution measurements.  The '4 Hz comb' was actually just the spectra plotting script switching between the light orange and dark orange spectra in Plot 2.  The light orange spectrum is taken at 1 Hz resolution, while the dark orange is taken at 4 Hz. The light orange, at higher resolution, should be the only spectrum reported at low frequency.  But if you leave in the low frequency dark orange points, the higher noise floor emerges as a 4 Hz frequency comb.  This is why hierarchical chucking is a very important process.

 

Attachment 1: NorthLowerPeriscopeMirror_2017-11-13_16.48.15.jpg  1.611 MB  Uploaded Tue Nov 14 13:47:22 2017  | Hide | Hide all
NorthLowerPeriscopeMirror_2017-11-13_16.48.15.jpg
Attachment 2: SouthLowerPeriscopeMirror_2017-11-13_16.49.35.jpg  1.664 MB  Uploaded Tue Nov 14 13:47:22 2017  | Hide | Hide all
SouthLowerPeriscopeMirror_2017-11-13_16.49.35.jpg
ELOG V3.1.3-